OpenStudy (anonymous):

evaluate 5^28 x 5^-32 x 5^-4

OpenStudy (anonymous):

the x is multiplication

OpenStudy (whpalmer4):

Is that supposed to be \[5^{28}*5^{-32}*5^{-4}\]

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

hint: \[x^n*x^m = x^{n+m}\]

OpenStudy (anonymous):

i cant read that. the numbers are blurry

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

When multiplying exponentials where the bases are identical, you keep the base the same and add the exponents.

OpenStudy (whpalmer4):

So for your problem, you keep the base (5), and add the 3 exponents: 28 + -32 + -4 = 28-36 =